A buffer with a pH of 4.33 contains 0.27 M of sodium benzoate and 0.20 M of benzoic acid. What is the concentration of [H+] in the solution after the addition of 0.056 mol of HCl to a final volume of 1.3 L? Assume that any contribution of HCl to the volume is negligible. [H+] = ?

--------

A buffer with a pH of 4.09 contains 0.17 M of sodium benzoate and 0.22 M of benzoic acid. What is the concentration of [H+] in the solution after the addition of 0.052 mol of HCl to a final volume of 1.3 L? Assume that any contribution of HCl to the volume is negligible.
Answer #1

1)

pH = 4.33

pH = pKa + log (sodium benzoate / benzoic acid)

4.33 = pKa + log [0.27 / 0.20]

pKa = 4.20

salt moles = 0.27 x 1.3 = 0.351

acid moles = 0.20 x 1.3 = 0.26

HCl moles = 0.056

pH =pKa + log [salt - HCl moless / acid + HCl moles ]

pH = 4.2 + log [0.351 -0.056 / 0.26 + 0.056 ]

pH = 4.17

[H+] = 10^-pH = 10^-4.17

[H+] =   6.76 x 10^-5 M

2)

pKa = 4.20

suppose initial volume is 1.3 litre

salt moles = 0.17 x 1.3 = 0.221

acid moles = 0.22 x 1.3 = 0.286

HCl moles = 0.052

pH =pKa + log [salt - HCl moless / acid + HCl moles ]

pH = 4.2 + log [0.221 -0.052 / 0.286 + 0.052 ]

pH = 3.90

[H+] = 1.26 x 10^-4 M